Instructions: Round your answers to the nearest dollar. a. Consumption housand b. Investment housand c. Government purchases thousand d. Net exports- thousand e. GDP thousandThe table below shows economic activity for a tiny country. Using the expenditure approach determine the following Total value (thousands of dollars) 700 100 Activity Families buy groceries Electronics c㎝pany sells HD projectors to households Personal trainer qives a Zumbaclass Custard stand sells pistachio ice cream Police department buys new cars Mayor leads creatio Elevator-construction company builds a new factor Local businessperson purchases corn from Mexico Sports-gear company sells hockey gloves to Canadian team Bike store sells five used carbon-fiber bikes 500 300 600 400 200 200 n of a new budget for education
Answer #1

Families buy groceries

Consumption

700

Electronics company buy

Consumption

100

Personal trainer

Consumption

8

Custard stand

Consumption

2

Police department

Government purchases

500

Mayor leads to

Investment

300

Elevator-construction

Investment

600

Local business

Import

400

Sports-gear

Exports

200

Bike store

Not included

200

a.Consumption

810

b.Investment

900

c.Government purchases

500

d.Net exports

Export-Import=-200

e.GDP

=C+I+G+NX=810+900+500-200=2010